OpenStudy (anonymous):

Use part 1 of the Fundamental Theorem of Calculus to find the derivative of the function.

OpenStudy (anonymous):

\[\int\limits_{2}^{1/x} \arctan t dt\]

OpenStudy (jamesj):

Let \[ F(t) = \int f(t) \ dt \] Then by the FTC, \( dF/dt = f(t) \). So use this with your problem. Let \( F(t) = \int \arctan t \ dt \). Then \[ \int_a^{1/x} \arctan t \ dt = F(1/x) - F(a) \] Now differentiate with respect to x.

OpenStudy (anonymous):

F'(1/x) (1/x^2)dx - F'(a) ? This is part 2 if I remember correctly

OpenStudy (jamesj):

not quite.

OpenStudy (jamesj):

Remember that F'(x) = arctan(x). Now use the chain rule carefully. Notice also that F(a) is a number, a constant. Hence it's derivative is .....

OpenStudy (anonymous):

I think I got the answer by \[h(x) = \int\limits_{2}^{1/x} \arctan t dt =- \arctan (1/x) *(1/x ^{2}) = -\arctan(1/x)/x ^{2}\]

OpenStudy (jamesj):

be more careful with your notation. The integral is not equal to the expressions on the right.

OpenStudy (jamesj):

but you have the right answer. Just write it carefully.

OpenStudy (anonymous):

I got the answer but still don't fully get what's going on

OpenStudy (jamesj):

Here's what is going on. Let g be the function defined by \[ g(x) = \int_a^{1/x} \arctan x \ dx \] We want to find dg/dx. That is what you've written down. Now there are a number of intermediate steps: let \( F(x) = \int \arctan x \ dx \), then \[ g(x) = F(1/x) - F(a) \] Therefore \[ \begin{align} \frac{dg}{dx} &= \frac{d\ }{dx} F(1/x) - \frac{d\ }{dx} F(a) \\ &= \frac{-1}{x^2} F'(1/x) - 0 \\ &= -\arctan(1/x)/x^2 \ \ \ \ \ \ \ \ \ \ \ \ \ \hbox{ by the FTC} \end{align} \]

OpenStudy (anonymous):

Thanks, the whole d/dx was throwing me for a loop.